OpenStudy (anonymous):

(cos Θ + cos Θ)^2 + (cos Θ + cos Θ)^2 answers are sin2 Θ 4 8 cos2 Θ

OpenStudy (anonymous):

SOMEONE PLEASE HELP

OpenStudy (astrophysics):

I just showed you this :\

OpenStudy (astrophysics):

Someone else isn't going to get a different answer...they will tell you your options are wrong as well, we got \[8 \cos^2(\theta)\]

OpenStudy (anonymous):

yes but i was hoping someone cou;d help me answer it maybe theres a different way

OpenStudy (astrophysics):

\[(\cos(\theta)+\cos(\theta)^2 = (2\cos \theta)^2 = 4\cos^2(\theta)\] since both sides are same we have \[4\cos^2(\theta)+4\cos^2(\theta) = 8\cos^2(\theta)\]

OpenStudy (astrophysics):

Your options don't make sense, either you just pick one at random or email your teacher

OpenStudy (anonymous):

okay i think im going to go with 8

OpenStudy (astrophysics):

Post the full question

OpenStudy (astrophysics):

Maybe you missed a detail

OpenStudy (astrophysics):

Including everything

OpenStudy (astrophysics):

Yeah ok the options are wrong, go either 8 or cos^2theta I guess :\ make sure to tell your teacher

OpenStudy (anonymous):

tahts all it is

OpenStudy (anonymous):

do you mind helping w another

OpenStudy (astrophysics):

sure

OpenStudy (anonymous):

what is the sine value of 5pi/3

OpenStudy (anonymous):

i know its equal to 300 degrees

OpenStudy (anonymous):

?

OpenStudy (astrophysics):

\[\sin(5 \pi/3) \] ok so we have this, the reference angle here is pi/3 what's the ratio of sin(theta)?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

im sorry im really bad at math i just know tis equal to 300 degrees

OpenStudy (astrophysics):

SOH CAH TOA

OpenStudy (astrophysics):

that ring a bell?

OpenStudy (astrophysics):

\[\sin( \theta) = \frac{ opposite }{ hypotenuse } = \frac{ x }{ r }\]

OpenStudy (anonymous):

oh yeah

OpenStudy (astrophysics):

 this is the pi/3 triangle, so far so good?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Ok since we know it's 300 degrees we know it's in the fourth quadrant of unit circle right. So\[y=-\sqrt{3}\]and\[r = 2\]\[\frac{ y }{ r } = \frac{ - \sqrt{3} }{ 2 }\] sorry I put x as opposite when it should be y/r and not x/r.

OpenStudy (anonymous):

okay

OpenStudy (astrophysics):

So that's our exact value

OpenStudy (anonymous):

thank you so much

OpenStudy (astrophysics):

Np